# Neighborhood Feature Pooling for Remote Sensing Image Classification

### v1

@@ 4.4 Similarity Measures
A total of 18 similarity metrics are evaluated within the NFP framework using ResNet-18 on the UC Merced dataset to identify the most effective function for capturing local feature relationships. As shown in Tab. 2, cosine similarity achieves the highest classification accuracy ($98.86\pm 0.58$), emphasizing the importance of angular alignment between feature vectors. This metric captures directional consistency while being invariant to magnitude, making it particularly suitable for remote sensing scenarios with variable illumination and intensity.
Several other metrics, including dot product ($98.25\pm 0.62$), Hellinger ($98.25\pm 0.56$), and Earth Mover’s Distance ($98.49\pm 0.59$), also show strong performance, suggesting that angle and distribution-aware functions are effective. In contrast, distance-based measures such as root mean square error (RMSE) ($98.02\pm 0.56$), L1 norm ($98.41\pm 0.40$), and L2 norm ($98.25\pm 0.68$) result in slightly lower accuracy, indicating that absolute deviations are less informative in texture-based classification tasks. Overall, these results highlight cosine similarity as the most appropriate choice for guiding NFP, due to its ability to preserve fine-grained spatial structure in local neighborhoods.
(a) Airplane
(b) GAP
(c) NFP (Ours)
(d) DeepTEN
(e) Fractal
(f) Lacunarity
(g) RADAM
Figure 4: Grad-CAM [23] visualizations for an “Airplane” sample from the UC Merced dataset [30] using MobileNetV3 [10] with different pooling strategies: (a) Original image, (b) GAP (Baseline), (c) NFP (Ours), (d) DeepTEN, (e) Fractal, (f) Lacunarity, and (g) RADAM. Brighter regions indicate higher model attention. NFP consistently yields more focused and semantically meaningful activations, localizing key object regions (e.g., airplane body and wings) better than GAP and other texture pooling methods. These results demonstrate NFP’s superior ability to preserve fine-grained, class-relevant spatial structure.
4.5 Layer Placement
Network Stage Channels Acc. $\pm$ Std (%) Params
@@ 4.5 Layer Placement
Table 3: Classification accuracy, number of channels before NFP (cosine similarity), and number of trainable parameters for different insertion stages of MobileNetV3 [10] on UC Merced [30]. Channels are from features_only output [26]. Best average result in bold.
@@
Figure 3 demonstrates how NFP modifies early feature representations compared to the baseline backbone. For the “intersection” scene, standard activations highlight edges and coarse textures, whereas NFP emphasizes neighborhood-consistent patterns such as road grids and parcel boundaries. This produces sharper, more spatially coherent responses, indicating that NFP reinforces structural regularities while reducing noisy activations. Such enhanced sensitivity to fine-grained arrangements provides qualitative support for the improvements observed in Tab. 1.
Figure 5: t-SNE visualizations of penultimate-layer features extracted from models trained on GTOS-Mobile (31 classes) with different pooling methods: GAP, NFP (Ours), DeepTEN, Fractal, Lacunarity, and RADAM. Each point represents a test image, colored by ground-truth class. The Silhouette Score (computed on the original feature space) is shown beneath each panel, quantifying the compactness and separability of class clusters. NFP achieves the highest Silhouette Score, indicating its superior ability to produce discriminative and well-separated feature embeddings for texture classification. All runs used a shared random seed for fair comparison.
4.6 Explainable AI analysis
To better understand the impact of different pooling strategies on model interpretability, Grad-CAM [23] is employed to visualize class activation maps for models trained with GAP (baseline), NFP, DeepTEN, Fractal, Lacunarity, and RADAM, as summarized in Fig. 4. Including GAP as a baseline enables direct comparison with the widely used global average pooling approach. The GAP model produced attention maps that highlighted regions that are only loosely aligned with the object of interest (airplane in this example). In contrast, NFP yields more semantically relevant activations, accurately localizing critical object parts (such as the airplane fuselage and wings). DeepTEN also sharpens attention relative to GAP, but remains more diffuse and less object-centered than NFP. The difference is particularly evident when comparing the concentrated attention in NFP to the more diffuse responses of GAP and RADAM, as well as the less object-aligned focus seen in Fractal and Lacunarity pooling. These results demonstrate that NFP not only improves classification accuracy but also enhances the interpretability of model predictions by producing more meaningful and focused visual explanations.
4.7 t-SNE Visualization Analysis
To evaluate the discriminative power of the learned feature embeddings, penultimate-layer features from each model are visualized using t-distributed Stochastic Neighbor Embedding (t-SNE) [24] on the GTOS-Mobile dataset.  Figure 5 presents two-dimensional projections for models utilizing GAP, Fractal, Lacunarity, RADAM, and the proposed NFP. Each point in the scatter plots corresponds to a test sample and different colors correspond to the true class. All visualizations are generated with identical t-SNE hyperparameters and random seed to ensure fair comparisons between models.
In addition to qualitative visual analysis, the Silhouette Score [21] for each method is reported beneath each plot. The Silhouette Score quantifies the compactness and separability of clusters in the original feature space, with higher values reflecting more well-defined and better-separated class clusters. As summarized in Fig. 5, NFP achieves the highest Silhouette Score ($0.4497$), indicating the most compact and well-separated clusters among all methods. This result aligns with the visual impression and shows that NFP produces more clearly defined and less overlapping clusters compared to GAP, Fractal, Lacunarity, and RADAM. In contrast, RADAM and GAP show lower Silhouette Scores ($0.3011$ and $0.3214$, respectively), consistent with their more diffuse, overlapping class distributions. In summary, both the qualitative t-SNE visualizations and the quantitative Silhouette Scores demonstrate that NFP enables more discriminative and structured feature embeddings, supporting its superior classification performance on the GTOS-Mobile dataset.
@@ 5 Conclusion
This work introduces NFP, a method designed to enhance texture-aware classification in remote sensing images. By explicitly modeling local similarity relationships within the feature space, NFP enables convolutional backbones to preserve fine-grained spatial structure alongside global semantic information. Extensive experiments across five public datasets and three backbone architectures demonstrate that NFP consistently matches or surpasses state-of-the-art pooling methods, including GAP, Lacunarity, Fractal, and RADAM, with up to 6.83% accuracy improvements on challenging benchmarks such as GTOS-Mobile. Importantly, these gains are achieved with minimal additional parameter cost.
Further analyses of similarity metrics, layer placement, and interpretability (e.g., Grad-CAM and t-SNE) confirm that NFP produces more discriminative and semantically meaningful feature embeddings compared to prior methods. Although the largest gains are observed on CNN-based models, NFP also delivers competitive results on transformer-based backbones. Future work include incorporating learnable weighting, radial or multi-scale neighborhoods, parameter-efficient transfer learning, extending NFP toward domain adaptation and lightweight deployment in diverse vision tasks (e.g. object detection and segmentation).
@@ Acknowledgment
This material is based upon work supported by Texas A&M AgriLife. Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing. LA-UR-25-30363.

### v3

@@ 4.4 Ablation: Similarity Measures
We evaluate 18 similarity measures within the NFP framework using ResNet-18 on the EuroSAT dataset to identify the most effective metric for modeling local feature relationships. As shown in Tab. 3, cosine similarity achieves the best classification performance ($98.52\pm 0.20$%), suggesting that angular alignment between local feature vectors is particularly informative for neighborhood encoding. Cosine similarity emphasizes directional agreement while being less sensitive to differences in feature magnitude, which is beneficial in multispectral remote sensing scenes with variations in illumination and contrast.
Several alternative measures also perform competitively, including Earth Mover’s Distance (EMD) ($97.09\pm 0.18$%), Squared Chord ($97.02\pm 0.17$%), and L2 norm ($96.96\pm 0.29$%). In contrast, performance is notably lower for metrics such as Canberra ($79.14\pm 1.02$%) and Chi-Squared Type 2 ($74.35\pm 7.66$%), indicating weaker suitability for local neighborhood comparison in this setting. Notably, Sharpened Cosine (SCS) exhibits substantially higher variance ($\pm 10.99$), suggesting unstable performance across runs. Overall, these results support cosine similarity as the default choice for NFP on EuroSAT, providing a strong balance between robustness and discriminative power for local neighborhood modeling.
4.5 Ablation: Neighborhood Dilation
To examine the effect of neighborhood sampling distance in NFP, a dilation-based ablation study is conducted on the RESISC45 dataset using a ResNet18 backbone with dilation factors of 1, 3, 5, 7, 9, 11, 13, and 15. Larger neighborhoods could be used to capture broader spatial context at the expense of reduced boundary precision. The neighborhood radius is fixed at $r=1$, and only the sampling distance is varied. As reported in Tab. 4, including interpolated intermediate values, performance differences across dilation settings remain minor, indicating that NFP is robust to changes in neighborhood spacing. A dilation of 3 achieves the highest average test accuracy, whereas larger dilation factors do not offer significant differences in performance.
Similarity measure Accuracy $\pm$ 1 Std (%)
Cosine 98.52 $\pm$ 0.20
Earth Mover’s Distance (EMD) 97.09 $\pm$ 0.18
Squared Chord 97.02 $\pm$ 0.17
L2 Norm 96.96 $\pm$ 0.29
Smith Similarity 96.96 $\pm$ 0.17
Chi-Squared Type 1 96.94 $\pm$ 0.05
Geman-McClure 96.88 $\pm$ 0.20
Root Mean Square Error (RMSE) 96.88 $\pm$ 0.17
Scaled Dot Product 96.83 $\pm$ 0.24
GFC 96.77 $\pm$ 0.12
Pearson Correlation 96.75 $\pm$ 0.35
Hellinger 96.68 $\pm$ 0.30
Jeffrey Divergence 96.23 $\pm$ 0.23
Sharpened Cosine (SCS) 87.67 $\pm$ 10.99
Canberra 79.14 $\pm$ 1.02
Chi-Squared Type 2 74.35 $\pm$ 7.66
L1 Norm 96.96 $\pm$ 0.29
Dot Product 96.83 $\pm$ 0.24
Table 3: Comparison of 18 similarity measures for NFP using ResNet-18 on the EuroSAT dataset. Best average result is shown in bold.
Dilation Factor Accuracy $\pm$ 1 Std (%)
1 94.02 $\pm$ 0.25
3 94.42 $\pm$ 0.10
5 94.16 $\pm$ 0.05
7 94.28 $\pm$ 0.07
9 94.11 $\pm$ 0.09
11 94.22 $\pm$ 0.06
13 94.08 $\pm$ 0.11
15 94.23 $\pm$ 0.04
Table 4: Effect of neighborhood dilation on NFP performance using ResNet18 on the RESISC45 dataset. Best average result is shown in bold. Intermediate values are interpolated for trend visualization.
4.6 Ablation: Layer Placement
Network Stage Channels Acc. $\pm$ 1 Std (%) Params
@@ 4.6 Ablation: Layer Placement
Table 5: Classification accuracy, number of channels before NFP (cosine similarity), and number of trainable parameters for different insertion stages of MobileNetV3 [9] on UC Merced [29]. Channels are from features_only output [27]. Best average result is shown in bold.
@@
4.7 Ablation: Initialization Strategy
EuroSAT has 13 spectral bands, so the first layer of the network must be modified to handle more input channels than the standard RGB case. The results in Tab. 6 show that mean channel-wise initialization gives the best accuracy for both GAP and the proposed NFP. This suggests that averaging the pretrained RGB weights provides a better starting point for learning from multispectral data. In contrast, random initialization removes useful pretrained information and leads to worse performance, while weight replication and the $1{\times}1$ adapter are less effective. In addition, the $1{\times}1$ adapter adds extra parameters, whereas mean initialization keeps the model lightweight. For these reasons, mean initialization is used in all other experiments.
4.8 Ablation: Fusion Strategy
Init. Strategy GAP (%) NFP (%)
Weight Replication $97.85\pm 0.25$ $98.05\pm 0.22$
Mean Init. $\mathbf{98.27\pm 0.19}$ $\mathbf{98.52\pm 0.20}$
$1{\times}1$ Adapter $97.42\pm 0.28$ $97.91\pm 0.26$
Random Init. $94.63\pm 0.83$ $94.80\pm 0.78$
Table 6: Effect of different 13-channel initialization (init) strategies on EuroSAT. Each entry reports average accuracy $\pm$ 1 standard deviation across multiple runs for both GAP and NFP. Mean initialization corresponds to the final configuration and yields the highest performance.
The proposed design combines NFP and GAP via element-wise product, allowing local neighborhood similarity responses to rescale global activations. To assess whether more expressive interactions are necessary, three alternative fusion mechanisms were evaluated on UC Merced with a ResNet18 backbone: (i) simple concatenation, (ii) concatenation followed by a two-layer MLP, and (iii) Squeeze-and-Excitation (SE) [10] style gated weighting, where a scalar weight $\alpha$ balances GAP and NFP features. As summarized in Tab. 7, the element-wise product achieves the highest accuracy while using the second fewest parameters. This suggests that the element-wise product preserves complementary information from both NFP and GAP, which creates a more aligned and interpretable interaction between the two branches, while avoiding unnecessary parameter growth. In contrast, the MLP and SE-style gating increase the model size from 11.23M to 12.02M and 12.28M parameters, respectively, without providing a proportional improvement in performance.
Fusion type Params (M) Acc. $\pm$ Std (%)
Concatenation only 11.19 96.35 ± 0.15
Concatenation + MLP 12.02 97.05 ± 0.13
Gated weighting (SE-style) 12.28 98.12 ± 0.25
Element-wise product (Ours) 11.23 98.86 ± 0.58
Table 7: Comparison of fusion strategies for combining NFP and GAP on the UC Merced dataset (ResNet18 backbone). Parameter count is given in millions. Best average result is shown in bold.
@@ 5 Conclusion
This work introduces NFP, a method designed to enhance texture-aware classification in remote sensing images. By explicitly modeling local similarity relationships within the feature space, NFP enables convolutional backbones to preserve fine-grained spatial structure alongside global semantic information. Extensive experiments across three public remote-sensing datasets and three backbone architectures demonstrate that NFP is competitive with, and often surpasses, state-of-the-art texture pooling methods, including GAP, Lacunarity, Fractal, and RADAM, as well as DeepTEN, while keeping the additional parameter cost minimal and preserving the lightweight footprint of the underlying backbones.
Further analysis of similarity metrics, layer placement, neighborhood dilation, and visualizations (e.g., t-SNE) confirm that NFP produces more discriminative and semantically meaningful feature embeddings compared to prior methods. A moderate dilation factor was found to slightly improve performance without increasing parameters, indicating that limited expansion of spatial context can be beneficial. These benefits are most pronounced on compact CNN backbones such as ResNet18, MobileNetV3, and ConvNeXt-Atto. Future work includes exploring learnable weighting schemes for combining GAP and NFP, more advanced multi-scale or adaptive neighborhood formulations (e.g., radial windows), as well as extending NFP toward domain adaptation and lightweight deployment in diverse vision tasks (e.g., object detection and segmentation).
@@ Acknowledgment
Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing. LA-UR-2530363.